There Are No New Homometric Golomb Ruler Pairs with 12 Marks or Less

Eric Postpischil Peter Gilbert

Abstract

We describe an exhaustive search that proves the nonexistence of homometric Golomb ruler pairs with 7 to 12 marks. We also give a two-dimensional geometric model for the known family of six-mark homometric ruler pairs.